OpenStudy (anonymous):

-5 - 8p = -3p A. p = -25 B. p = 25 C. p = 1 D. p = -1

OpenStudy (anonymous):

D

OpenStudy (anonymous):

-5-8p=-3p -5=-3p+8p -5=5p p=-1

OpenStudy (anonymous):

So D